# Orthogonal Low-rank Adaptation in Lie Groups for Continual Learning of Large Language Models

### v1

@@ Abstract
Large language models (LLMs) are prone to catastrophic forgetting in sequential multi-task settings. Parameter regularization methods such as O-LoRA and N-LoRA alleviate task interference by enforcing low-rank subspace orthogonality, but they overlook the fact that conventional additive fine-tuning disrupts the intrinsic geometric structure of LLM parameters, limiting performance. Our key insight is that the parameter space of LLMs possesses a geometric structure, which must be preserved in addition to enforcing orthogonality. Based on this, we propose Orthogonal Low-rank Adaptation in Lie Groups (OLieRA), which introduces Lie group theory into LLM fine-tuning: leveraging multiplicative updates to preserve parameter geometry while applying orthogonality constraints to task subspaces. Experiments demonstrate that OLieRA achieves state-of-the-art results on the Standard CL benchmark and remains among the top-performing methods in the Large Number of Tasks setting. Moreover, OLieRA inherits the core advantages of O-LoRA—no need for replay data and task-ID-free inference—establishing a new paradigm for continual learning in LLMs.
@@
Kefan Cao  and Shuaicheng Wu Yingcai Honor College, University of Electronic Science and Technology of China
@@ 1 Introduction
Enabling lifelong continual learning in large language models (LLMs) is crucial for practical AI, yet they are prone to catastrophic forgetting across sequential tasks. To address the catastrophic forgetting issue McCloskey and Cohen (1989) faced by large language models (LLMs) in sequential multi-task scenarios, the research community has proposed continual learning as a solution—aiming to incrementally accumulate knowledge in non-stationary task streams while preventing the degradation of performance on previously learned tasks.
@@
Figure 1: Illustration of the update mechanisms of OLieRA and O-LoRA. While O-LoRA updates parameters orthogonally to previously updated dimensions, it overlooks the intrinsic parameter structure. In contrast, OLieRA incorporates a Lie group constraint in addition to orthogonality, thereby preserving the original parameter structure.
Existing continual learning approaches can be broadly categorized into three classes, each with notable limitations: (1)Replay-based methods  de Masson d’Autume et al. (2019): These methods store historical task data in a memory buffer and allow the model to revisit these samples during new task training. However, they require continuous storage of potentially sensitive user data, raising privacy concerns. Additionally, updating all model parameters leads to high computational costs. (2)Regularization-based methods Kirkpatrick et al. (2017): These approaches introduce an additional regularization term in the loss function to penalize significant changes to important weights, thereby preserving knowledge from previous tasks. Nevertheless, most methods in this category suffer from degraded generalization in long task sequences and struggle to adapt to continuous multi-task learning settings. (3)Architecture-based methods Razdaibiedina et al. (2023): These techniques dynamically expand model capacity or isolate task-specific parameters to minimize interference. However, they essentially construct separate sub-models for different tasks, limiting their generalization to unseen tasks and reducing flexibility and universality.
Our work falls within the regularization-based paradigm and builds upon Orthogonal Low-Rank Adaptation (O-LoRA) Wang et al. (2023a). Instead of updating parameters across a shared vector space for all tasks, O-LoRA performs gradient updates in directions orthogonal to the subspaces of previous tasks. This strategy has been shown to effectively mitigate catastrophic forgetting by avoiding interference with the loss landscapes of historical tasks.
In this paper, we propose Orthogonal Low-Rank Adaptation under Lie Groups (OLieRA). Unlike O-LoRA, which maintains orthogonality only within a specific low-rank subspace, OLieRA enforces orthogonality constraints directly across entire task subspaces. Moreover, it leverages Lie group theory to handle task subspaces. Compared to O-LoRA, OLieRA introduces only minimal additional computational overhead while achieving improved performance and enhanced theoretical interpretability. Figure 1 illustrates the conceptual differences between the update mechanisms of OLieRA and O-LoRA.
Thus, OLieRA retains all three key advantages of O-LoRA:
  • a.
    Data Privacy-Friendly: Eliminates the need to store user data for replay, thereby resolving data privacy concerns at the source;
  • b.
    Parameter-Efficient: Introduces only a small number of additional parameters, enabling efficient learning of new tasks without compromising performance on previous ones;
  • c.
    Generalization-Friendly: Does not require task identity during testing, naturally aligns with instruction tuning paradigms, and effectively preserves the generalization ability of LLMs to unseen tasks.
Furthermore, OLieRA achieves comprehensive performance improvements, surpassing all existing methods in the orthogonal low-rank adaptation direction.
@@
    We propose OLieRA, a parameter regularization-based continual learning framework grounded in Lie group theory. It significantly outperforms existing regularization-based methods such as O-LoRA and N-LoRA, achieving state-of-the-art (SOTA) results on the Standard CL Benchmark;
@@
    We apply Lie group theory to the fine-tuning process of LLMs, providing a geometric perspective for parameter updates that better preserves the intrinsic structure of the model’s representation space. Extensive analytical experiments validate the effectiveness and theoretical soundness of the proposed approach.
@@ 2.3 O-LoRA
O-LoRA Wang et al. (2023a) is a framework for continual learning in language models. It first integrates human expertise and enhances generalization through instruction tuning. Then, for each incoming task, it incrementally learns a new LoRA module while enforcing orthogonality between the current task’s LoRA parameters and those of previous tasks.
We further elaborate on the notion of orthogonality in O-LoRA: Orthogonality in O-LoRA is enforced between the low-rank subspaces spanned by the LoRA parameters of different tasks.
Specifically, for each task $t$, the model learns a set of LoRA parameters $\{A_{t},B_{t}\}$, where $B_{t}\in\mathbb{R}^{d\times r_{t}}$ ($d$ is the model dimension, $r_{t}$ is the low rank). The parameter update subspace $\mathcal{U}_{t}$ for task $t$ is approximated as the span of the column vectors of $B_{t}$:
$\mathcal{U}_{t}=\text{span}\{\mathbf{b}_{t}^{1},\mathbf{b}_{t}^{2},...,\mathbf{b}_{t}^{r_{t}}\}$ (2)
where $\mathbf{b}_{t}^{i}$ denotes the $i$-th column vector of $B_{t}$.
To prevent new tasks from interfering with past ones, O-LoRA requires the current task’s subspace $\mathcal{U}_{t}$ to be orthogonal to all past task subspaces $\mathcal{U}_{i}$ ($i<t$). This orthogonality is enforced by constraining $B_{t}$ such that:
$O_{i,t}=B_{i}^{\top}B_{t}=\mathbf{0}$ (3)
That is, the transpose of $B_{i}$ (for $i=1,\dots,t-1$) multiplied by $B_{t}$ yields a zero matrix, ensuring that the inner product of any vectors from the two subspaces is zero ($\langle\mathbf{u},\mathbf{w}\rangle=0$, $\forall\mathbf{u}\in\mathcal{U}_{i},\mathbf{w}\in\mathcal{U}_{t}$).
During training, an orthogonality loss term is added to the objective:
$L_{\text{orth}}(B_{i},B_{t})=\sum_{j,k}\left|O_{i,t}[j,k]\right|^{2}$
where $O_{i,t}[j,k]$ denotes the element at the $j$-th row and $k$-th column of $O_{i,t}$. This reinforces the orthogonal constraint, minimizes inter-task interference, and mitigates catastrophic forgetting.
@@ 2.5 LieRA
LieRA is a recently proposed parameter-efficient fine-tuning method. Unlike standard LoRA, Si et al. (2025) treats parameters as elements of a Lie group and models updates as perturbations in the Lie algebra, preserving structure via the exponential map. Instead of the additive update $W+\Delta W$ used in standard LoRA, LieRA employs a multiplicative update:
@@
where $\exp(\Delta W)$ denotes the exponential mapping. This approach has demonstrated superior performance across various domains including computer vision (image classification, object detection/segmentation, image generation) and natural language processing (commonsense reasoning, language understanding). This fine-tuning method helps preserve the intrinsic structure of high-dimensional parameters (e.g., spatial locality in convolutional kernels), enhancing feature capture capability. The Lie group-Lie algebra framework ensures smooth and consistent updates, improving the model’s flexibility in adapting to new tasks. It retains the efficiency of low-rank adaptation while enhancing expressiveness through Hadamard product-based updates. Furthermore, it offers a geometric perspective on parameter updates.
@@ 3.1 Introducing Lie Groups
To interpret parameter updates using Lie groups, we first need to construct a Lie group and attempt to describe the parameters within its framework. We consider collecting all possible instances of the parameters $W$ into a set. It then suffices to define a well-defined operation on this set such that it satisfies the group structure, forms an Abelian group, and is a smooth manifold—thereby constituting a Lie group.
Define the set $G=\{W\in\mathbb{R}^{b_{1}\times b_{2}\times\dots\times b_{k}}\mid W[a_{1},a_{2},\dots,a_{k}]\neq 0,\forall(a_{1},a_{2},\dots,a_{k})\in([1,b_{1}],[1,b_{2}],\dots,[1,b_{k}])\}$. For practical reasons and operational convenience, we assume all elements are non-zero. A suitable approach is to use the Hadamard product $\odot$ (element-wise multiplication) as the operation on this set. This is well-defined and causes all $W$ to form a group, as it effectively reduces the operation to one-dimensional non-zero real multiplication, which itself is well-defined and forms a group. Furthermore, since multiplication on one-dimensional non-zero real numbers is commutative and a smooth manifold, $(G,\odot)$ constitutes an Abelian group and a smooth manifold—i.e., a Lie group.
A rigorous verification is as follows. Note that:
(1) The element $I$ with all entries equal to 1 serves as the identity element.
(2) Since every entry in any $W$ is non-zero, the Hadamard product of any two $W$ matrices also has all non-zero entries, satisfying closure.
(3) Associativity follows from the associativity of one-dimensional real multiplication.
(4) Commutativity follows from the commutativity of one-dimensional real multiplication.
(5) Every element has an inverse (element-wise reciprocal), which exists due to the non-zero constraint.
Thus, $(G,\odot)$ is an Abelian group.
@@ 3.2 Introducing Lie Algebras
If the Lie group $(G,\odot)$ is viewed as a surface, then $W$ belonging to this Lie group represents a point on that surface. The Lie algebra is the tangent plane at that point relative to the surface. Movement on the tangent plane is simpler because it is linear. Formally, the Lie algebra is a linear space. Treating $W$ as a real vector, we can perform scalar multiplication and addition, forming a linear space denoted $\mathfrak{g}$.
Why introduce the Lie algebra? Because direct updates within the Lie group are cumbersome. Normally, to update, we would have: $W_{\text{new}}=W_{\text{old}}\odot\Delta W$, where $W_{\text{new}},W_{\text{old}},\Delta W\in G$. However, during optimization, we first compute the "required $W_{\text{new}}$" based on the loss function, and then invert to find the "amount to ’add’ $\Delta W$". This necessitates using the inverse element to solve for $\Delta W$: By "left-multiplying" both sides by the inverse of $W_{\text{old}}$, denoted $W_{\text{old}}^{-1}$ (a property of groups: every element has an inverse satisfying $W_{\text{old}}\odot W_{\text{old}}^{-1}=I$), we obtain:
$\Delta W=W_{\text{old}}^{-1}\odot W_{\text{new}}$
Therefore, it is preferable to first perform updates in the Lie algebra $\mathfrak{g}$ and then map back to the Lie group via a specific mechanism. This returns us to the familiar tensor computations (addition and scalar multiplication) used in frameworks like PyTorch.
@@ 3.3 Improved Incremental Update Mechanism
For any $W\in G$, we do not directly modify $W$ itself. Instead, we construct a small perturbation tensor $\Delta W$ in the corresponding Lie algebra $\mathfrak{g}$. $\Delta W$ is an ordinary tensor fully compatible with deep learning frameworks: gradients can be computed directly via .grad, linear updates can be performed using .add_() or .mul_(), and it can be adapted via reshape to existing fine-tuning methods designed for 2D tensors (e.g., LoRA), requiring no modifications to the optimization pipeline.
Figure 3: Illustration of the difference between the improved incremental update mechanism and the previous approach.
After performing the linear update on $\Delta W$ in the Lie algebra $\mathfrak{g}$, it must be mapped "smoothly" back to the Lie group $G$. Directly updating via the Lie group multiplication $W\odot\Delta W$ is problematic because $\Delta W$ might not satisfy the constraints of $G$ (e.g., all entries non-zero), potentially moving the result outside the valid parameter space. The exponential map $\exp:\mathfrak{g}\to G$ is the crucial "bridge" connecting the Lie algebra to the Lie group. Its key roles are threefold:
  1. 1.
    Local Homeomorphism: For infinitesimal perturbations $\Delta W$, $\exp(\Delta W)$ maintains a local continuous correspondence with $\Delta W$, ensuring the result always lies within the valid parameter space of $G$ and avoids parameter boundary issues (e.g., elements becoming zero);
  2. 2.
    Structure Preservation: The result of $\exp(\Delta W)$ necessarily belongs to $G$, and the mapping is implemented via element-wise exponential scaling—only adjusting the numerical proportions of the weights without disrupting the tensor dimensions or topological structure of the parameter space;
  3. 3.
    Geometric Consistency: It guides weight updates along the natural gradient direction of the Lie group manifold, avoiding the issue of "update direction decoupling from manifold geometry" inherent in ordinary linear addition (e.g., $W_{\text{new}}=W_{\text{old}}+\Delta W$).
Based on this, the weight update formula can be expressed as:
$W\mapsto W\odot\exp(\Delta W)$ (6)
This formulation ensures the legality of the group operation via the Lie group multiplication $\odot$ (Hadamard product), while leveraging the linear properties of the Lie algebra to simplify the optimization of $\Delta W$, achieving a balance between "structural constraints" and "optimization efficiency". Figure 3 visually illustrates the difference before and after this improvement.
3.4 First-Order Taylor Approximation of the Exponential Map
@@ 3.6 Low-Rank Decomposition
$\begin{split}W\odot\exp(\Delta W)&\approx W\odot(I+\Delta W)\\ &=W+W\odot\Delta W\\ &=W+W\odot(BA)\end{split}$ (11)
@@ 5.1 Interpretation via Lie Group Introduction
Inspired by Si et al. (2025), we can interpret the pre-trained parameters $W$ as elements of a Lie group. The fine-tuning adjustment $\Delta W$ applied to these parameters can then be viewed as an infinitesimal perturbation of this element. The elements of the corresponding Lie algebra constitute a specific collection of such infinitesimal perturbations (i.e., perturbations on the tangent plane). Crucially, infinitesimal perturbations from the Lie algebra can be mapped back to the Lie group via structure-preserving maps, such as the exponential map, ensuring the perturbed element remains within the group. This is the fundamental rationale behind Lie algebras: direct additive changes to Lie group elements do not guarantee the result stays within the group. The Lie algebra formalism, coupled with a structure-preserving mapping like the exponential map, allows us to constrain the parameter update $\Delta W$ to be a valid infinitesimal perturbation (satisfying Lie algebra conditions), thereby achieving the desired effect of preserving the Lie group structure.
A prerequisite for this approach is that $W$ itself possesses a structure worth preserving. Si et al. (2025) demonstrated that convolutional kernels indeed benefit from such structure-preserving constraints during fine-tuning. We posit that Large Language Models (LLMs) also contain structures that need preservation. While the structure in LLMs differs from the spatial locality crucial for convolutional kernels and is more challenging to interpret directly due to the model’s complexity, some explanations can be offered.
LLMs comprise diverse layers, such as the Q/K/V matrices in Transformer layers, Feed-Forward network weights, layer normalization parameters, etc. First, these modules exhibit internal structures, much like trained convolutional kernels. Certain aspects of these structures must be preserved, as even minor alterations could catastrophically disrupt performance on previously learned tasks—the essence of catastrophic forgetting. Second, these modules are not isolated; they are intricately interconnected with complex, abstract relationships. Preserving these inter-modular relationships is another crucial aspect of the overall structure.
Based on this reasoning, we argue that employing Lie group and Lie algebra theory for parameter updates to preserve structure is necessary. Consequently, we use $W\odot\exp(\Delta W)$ for parameter updates instead of $W+\Delta W$. Furthermore, we can still apply low-rank decomposition to $\Delta W$ within $\exp(\Delta W)$, thus retaining the advantages associated with low-rank adaptation.

### v2

@@ Abstract
Large language models (LLMs) suffer from catastrophic forgetting in sequential multi-task learning. Existing parameter regularization methods (e.g., O-LoRA, N-LoRA) mitigate interference via low-rank subspace orthogonality, but additive updates distort the intrinsic geometry of model parameters. We propose OLieRA, a Lie group–based fine-tuning framework that preserves parameter geometry through multiplicative updates while enforcing orthogonality across task subspaces. OLieRA achieves state-of-the-art performance on the Standard CL benchmark and remains highly competitive under large task sequences. It further inherits the replay-free and task-ID–free inference properties of O-LoRA, establishing a principled paradigm for continual learning in LLMs.
@@
Kefan Cao  and Shuaicheng Wu Yingcai Honor College, University of Electronic Science and Technology of China kefancao@std.uestc.edu.cn
@@ 1 Introduction
Enabling lifelong continual learning in large language models (LLMs) is essential for real-world deployment, yet they remain highly susceptible to catastrophic forgetting, whereby acquiring new tasks degrades performance on previously learned ones (McCloskey and Cohen, 1989). This limitation is especially severe in practical settings, where task adaptation (e.g., sentiment analysis) can impair prior capabilities such as technical document summarization Kotha et al. (2024). Continual learning addresses this challenge by enabling incremental knowledge acquisition from nonstationary task streams while preserving performance on historical tasks Parisi et al. (2019).
@@
Figure 1: Illustration of the update mechanisms of OLieRA. While LoRA updates parameters, it overlooks the intrinsic parameter structure. In contrast, OLieRA incorporates a Lie group constraint in addition to orthogonality, thereby preserving the original parameter structure.
Existing continual learning methods fall into three paradigms, each with inherent constraints. Replay-based methods de Masson d’Autume et al. (2019); Lopez-Paz and Ranzato (2017) rely on storing past data, incurring privacy risks and prohibitive computational costs for large LLMs. Regularization-based methods Kirkpatrick et al. (2017) constrain parameter drift but degrade over long task sequences and struggle with ambiguous task boundaries PENG et al. (2024); Ke and Liu (2023). Architecture-based methods Razdaibiedina et al. (2023); Wang et al. (2023c) expand or isolate parameters to reduce interference, but fragment model capacity and limit transfer to unseen tasks.
Our work builds on the regularization paradigm, extending Orthogonal Low-Rank Adaptation (O-LoRA) Wang et al. (2023a). While O-LoRA mitigates interference via orthogonal gradient updates, its additive updates ($W\rightarrow W+\Delta W$) can distort the intrinsic geometry of LLM parameters Si et al. (2025). We propose Orthogonal Low-Rank Adaptation under Lie Groups (OLieRA), which enforces orthogonality across full task subspaces and preserves geometric structure through Lie group–based updates Si et al. (2025), with minimal additional overhead and improved theoretical interpretability (Fig. 1).
OLieRA retains the key advantages of O-LoRA: privacy-friendliness (no data replay), parameter efficiency, and task-agnostic generalization. It consistently outperforms prior orthogonal low-rank methods, including N-LoRA Yang et al. (2025).
@@
    We introduce OLieRA, a Lie group–based regularization framework that achieves state-of-the-art results on standard continual learning benchmarks, surpassing O-LoRA and N-LoRA;
@@
    We provide a geometric view of LLM fine-tuning, demonstrating that Lie group–constrained updates better preserve parameter structure, supported by extensive analytical experiments.
@@ 2.3 O-LoRA
O-LoRA Wang et al. (2023a) is a continual learning framework for language models that combines instruction tuning with task-specific low-rank adaptation. For each new task, it incrementally adds a LoRA module while enforcing orthogonality between the parameter subspaces of the current and previous tasks to reduce interference.
For task $t$, O-LoRA learns LoRA parameters $\{A_{t},B_{t}\}$ with $B_{t}\in\mathbb{R}^{d\times r_{t}}$. The task-specific update subspace is defined as
$\mathcal{U}_{t}=\mathrm{span}\{\mathbf{b}_{t}^{1},\ldots,\mathbf{b}_{t}^{r_{t}}\},$ (2)
where $\mathbf{b}_{t}^{i}$ denotes the $i$-th column of $B_{t}$. To prevent catastrophic forgetting, O-LoRA enforces orthogonality between subspaces by constraining
$B_{i}^{\top}B_{t}=\mathbf{0},\quad\forall i<t.$ (3)
This constraint is implemented via an orthogonality regularizer
$L_{\text{orth}}(B_{i},B_{t})=\|B_{i}^{\top}B_{t}\|_{F}^{2},$ (4)
which minimizes inter-task interference and preserves previously learned knowledge.
@@ 2.5 LieRA
LieRA is a recently proposed parameter-efficient fine-tuning method. Unlike standard LoRA, which applies additive updates, LieRA Si et al. (2025) models parameters as elements of a Lie group and performs updates in the associated Lie algebra, mapping them back via the exponential map. This yields a multiplicative update,
@@
that explicitly preserves parameter structure. LieRA has shown consistent gains across vision and language tasks, including image understanding and commonsense reasoning. By maintaining the intrinsic structure of high-dimensional parameters (e.g., spatial locality in convolutional kernels), the Lie group–Lie algebra framework enables smooth, consistent updates, improves adaptability to new tasks, and retains the efficiency of low-rank adaptation while offering a principled geometric view of parameter updates.
@@ 3.1 Introducing Lie Groups
Lie groups have been widely applied in machine learning for manifold optimization, enabling parameter updates that respect intrinsic geometric structure Leake and Vishnoi (2021). Representative applications include orthogonal parameterization Plumbley (2004) and modeling geometric transformations in computer vision Liu et al. (2023); Shutty and Wierzynski (2023). In this work, we extend Lie group methods to the continual learning setting. To model parameter updates via Lie theory, we embed the parameter tensor space into a Lie group. Let
$\begin{split}G=&\{W\in\mathbb{R}^{b_{1}\times\cdots\times b_{k}}\mid W[a_{1},\ldots,a_{k}]\neq 0\\ &\forall(a_{1},\ldots,a_{k})\},\end{split}$ (7)
and define the group operation as the Hadamard product $\odot$. Requiring all entries to be non-zero ensures well-defined inverses and simplifies the structure Varadarajan (2013).
Under $\odot$, $G$ is closed, associative, and commutative; the all-ones tensor serves as the identity, and each element admits an element-wise inverse. Since the operation is smooth, $(G,\odot)$ is an Abelian Lie group, isomorphic to a product of copies of $(\mathbb{R}^{\ast},\times)$.
@@ 3.2 Introducing Lie Algebras
Viewing the Lie group $(G,\odot)$ as a smooth manifold, a parameter set $W\in G$ corresponds to a point on this manifold, while the associated Lie algebra $\mathfrak{g}$ is the tangent space at $W$ (Fig. 1). The Lie algebra is a linear space, where updates are simpler due to the availability of addition and scalar multiplication.
The Lie algebra is introduced because direct optimization on the Lie group is inconvenient. In principle, group updates take the form $W_{\text{new}}=W_{\text{old}}\odot\Delta W$ with all variables in $G$. However, gradient-based optimization typically computes additive updates, requiring the inversion
$\Delta W=W_{\text{old}}^{-1}\odot W_{\text{new}},$ (8)
which complicates learning on the group. Instead, we perform updates in the Lie algebra $\mathfrak{g}$ and map them back to the Lie group via a suitable mapping (e.g., the exponential map), thereby recovering standard tensor operations compatible with frameworks such as PyTorch.
@@ 3.3 Improved Incremental Update Mechanism
Recent studies Marchetti et al. (2025); Modell et al. (2025) suggest that neural network parameters lie on a largely smooth manifold, which can be interpreted as a Lie group. Given $W\in G$, we introduce an infinitesimal perturbation $\Delta W\in\mathfrak{g}$, perform linear updates in the Lie algebra, and map the result back to the group via the exponential map $\exp:\mathfrak{g}\!\to\!G$. Direct group-wise updates (e.g., $W\odot\Delta W$) are ill-posed, as $\Delta W$ may violate the constraints of $G$. The exponential map thus provides a principled bridge from $\mathfrak{g}$ to $G$.
Specifically, $\exp(\Delta W)$ (i) ensures local validity by mapping infinitesimal perturbations to admissible group elements, (ii) preserves structure by remaining within $G$ through element-wise exponential scaling, and (iii) maintains geometric consistency by aligning updates with the manifold geometry rather than arbitrary linear directions. Consequently, we adopt the update
$W\rightarrow W\odot\exp(\Delta W),$ (9)
which combines structure-preserving group multiplication with efficient optimization in the linear Lie algebra, balancing geometric constraints and computational efficiency (Fig. 1).
3.4 Taylor Approximation of the Exponential Map
@@ 3.5 Low-Rank Decomposition
$\begin{split}W\odot\exp(\Delta W)&\approx W\odot(I+\Delta W)\\ &=W+W\odot(BA)\end{split}$ (14)
@@ 5.1 Interpretation via Lie Group Introduction
Inspired by Si et al. (2025), we view the pre-trained parameters $W$ as elements of a Lie group, with fine-tuning updates $\Delta W$ interpreted as infinitesimal perturbations. Such perturbations naturally reside in the associated Lie algebra, i.e., the tangent space of the group. Through structure-preserving mappings such as the exponential map, these infinitesimal updates can be mapped back to the Lie group, ensuring that the updated parameters remain within the group. This framework addresses a fundamental issue: naive additive updates generally do not preserve Lie group structure. By constraining $\Delta W$ to satisfy Lie algebra conditions and applying the exponential map, we obtain principled, structure-preserving parameter updates.
This approach requires that $W$ encodes meaningful structure. Prior work has shown that convolutional kernels benefit from structure-preserving fine-tuning Si et al. (2025). We argue that Large Language Models (LLMs) similarly exhibit structures that should be preserved. Although these structures differ from the spatial locality of convolutional kernels and are harder to interpret, they are nonetheless critical. LLMs consist of heterogeneous yet highly structured components (e.g., Q/K/V matrices, feed-forward layers, and normalization parameters), whose internal organization must be maintained to avoid catastrophic forgetting. Moreover, these components are tightly coupled, and preserving their interdependencies is essential for stable adaptation.
Motivated by these observations, we adopt Lie group–based parameter updates of the form $W\odot\exp(\Delta W)$ instead of $W+\Delta W$. Importantly, $\Delta W$ can still be parameterized via low-rank decomposition inside $\exp(\Delta W)$, thereby retaining the efficiency and flexibility of low-rank adaptation while enforcing structural consistency.